OpenStudy (anonymous):

Simplify. Express the product as a radical expression. (x^4) ^3/5

OpenStudy (anonymous):

\[(x ^{4}) ^{3/5}\]

jhonyy9 (jhonyy9):

so do you knoe the exponentts property ?

jhonyy9 (jhonyy9):

know the property of exponents ?

OpenStudy (anonymous):

not really

jhonyy9 (jhonyy9):

so (x^a)^b =x^(a*b) ok ?

OpenStudy (anonymous):

....ok ..

jhonyy9 (jhonyy9):

and if sqrtx =x^(1/2) so than your exercise how will be formed ?

OpenStudy (anonymous):

im not gunna lie.. im so confused and I have no idea what you are saying or doin

jhonyy9 (jhonyy9):

so there mean that when you have there a term with fractional exponent so than this sign that the denominator always will be the indice of radical ok ?

OpenStudy (anonymous):

.....

jhonyy9 (jhonyy9):

so if you multiplie 4 by 3/5 =12/5 than will result x^(12/5) what expressed in radical form so this will be radical indice 5 from x on exponent 12 ok ?

OpenStudy (anonymous):

alright...

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

ok I tried and I still aint got it